# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ Abstract
The identification of DNA-binding proteins (DBPs) is a critical task due to their significant impact on various biological activities. Understanding the mechanisms underlying protein-DNA interactions is essential for elucidating various life activities. In recent years, machine learning-based models have been prominently utilized for DBP prediction. In this paper, to predict DBPs, we propose a novel framework termed a multiview random vector functional link (MvRVFL) network, which fuses neural network architecture with multiview learning. The proposed MvRVFL model combines the benefits of late and early fusion, allowing for distinct regularization parameters across different views while leveraging a closed-form solution to determine unknown parameters efficiently. The primal objective function incorporates a coupling term aimed at minimizing a composite of errors stemming from all views. From each of the three protein views of the DBP datasets, we extract five features. These features are then fused together by incorporating a hidden feature during the model training process. The performance of the proposed MvRVFL model on the DBP dataset surpasses that of baseline models, demonstrating its superior effectiveness. Furthermore, we extend our assessment to the UCI, KEEL, AwA, and Corel5k datasets, to establish the practicality of the proposed models. The consistency error bound, the generalization error bound, and empirical findings, coupled with rigorous statistical analyses, confirm the superior generalization capabilities of the MvRVFL model compared to the baseline models.
@@
Keywords: Multiview learning, Support vector machine, Random vector functional link network, Extreme learning machine, DNA-binding protein.
1 Introduction
@@
The protein capable of interacting with DNA is known as a DNA-binding protein (DBP) (Jones et al., 1987). The DBP plays a crucial role in numerous vital biological processes, including transcriptional regulation, DNA replication and repair, cellular development, and chromatin organization (Ohlendorf et al., 1982). Therefore, precise prediction of DBPs is of considerable significance for proteome annotation as well as endeavors in synthetic biology. To identify DBPs, extensive work has been carried out in wet lab settings, including techniques such as X-ray crystallography (Chou et al., 2003), genetic analysis (Freeman et al., 1995), and chromatin immunoprecipitation on microarrays (Buck & Lieb, 2004). While results obtained from wet-lab methods are likely the most reliable, it’s important to note that this approach is also associated with significantly high time and labor costs. In contrast to wet-lab methods, computational approaches can substantially decrease resource requirements and expedite the identification of DBPs. Furthermore, in the postgenomic era, there’s an escalating demand for the development of efficient and swift computational methods for accurately identifying DBPs, thereby underscoring its increasing significance in the field of bioinformatics.
@@
Currently, computational methods are extensively employed for predicting DBPs (Qu et al., 2019). These methods are typically categorized into two groups: structure-based and sequence-based methods. Structure-based methods necessitate information obtained from the three-dimensional structure of the protein under examination. In the early stages, structure-based methods excelled in the domain of DBP prediction, exemplified by models like LBi-DBP (Zeng et al., 2024), StackDPP (Ahmed et al., 2024), DBPboost (Sun et al., 2024), and ULDNA (Zhu et al., 2024). Accurately representing a protein sequence as a vector is widely recognized as one of the sequence-based methods’ most essential and challenging tasks (Zhang & Liu, 2019). DNABinder (Ranjan et al., 2021) uses evolutionary information encoded in a position-specific scoring matrix (PSSM), generated via the PSI-BLAST multiple sequence alignment tool (Altschul et al., 1997a). A support vector machine (SVM) is employed for classification, using these features as input, marking the first instance of this approach being used to identify DNA-binding proteins (DBPs). In iDNA-Prot (Chou, 2011), pseudo amino acid composition (PseAAC) features (Chowdhury et al., 2017), extracted from protein sequences using the grey model, are amalgamated with random forest (RF) for the identification of DBPs. iDNAProt-ES (Chowdhury et al., 2017) utilizes evolutionary data like PSSM alongside structural information predicted by SPIDER2 (Yang et al., 2017) to characterize the features of a given protein. These features are trained to identify DBPs using SVM with a linear kernel. DBP datasets, derived from various sources or views, offer complementary information essential for improving task performance. Each view contributes unique perspectives and features that, when combined, enhance the overall understanding and efficiency of the model. Nevertheless, existing multiview learning approaches often necessitate complex architectures or extensive preprocessing to manage these diverse views effectively (Hu et al., 2023). The growing complexity of data from different sources poses significant challenges in extracting meaningful insights and achieving high performance. Multiview information fusion has gained widespread use in the field of bioinformatics in recent years (Zou et al., 2019; Zhou et al., 2022). The fuzzy kernel ridge regression model based on multi-view sequence features (FKRR-MVSF) (Zou et al., 2019) is proposed for identifying DNA-binding proteins. In FKRR-MVSF, the initial step is to extract multi-view sequence features from the protein sequences. Then, a multiple kernel learning (MKL) algorithm is used to combine these multiple features. In MSFBinder (Liu et al., 2018), a stacking framework is proposed to predict DBPs by combining features from multiple views. The multi-view hypergraph restricted kernel machines (MV-H-RKM) (Guan et al., 2022) model is proposed to extract multiple features from protein sequences. These features are connected via a common hidden feature, and multi-hypergraph regularization is applied to merge the multi-view features, maintaining structural consistency between the original and hidden features.
Despite the extensive use of hyperplane-based classifiers like SVMs and their variants in predicting DNA-binding proteins, a significant gap remains in exploring the potential of shallow neural networks for this task. Shallow neural networks, with their ability to learn complex data representations and capture non-linear relationships through hidden layers, present a compelling alternative to traditional machine learning models. However, conventional neural network models often face challenges like slow convergence, difficulties with local minima, and sensitivity to learning rates and initialization points, resulting in suboptimal outcomes. This challenge underscores the need for innovative yet simple approaches that can effectively manage diverse data types and leverage their full potential. At this juncture, the random vector functional link (RVFL) network emerges as a promising solution. Known for its simplicity and effectiveness across various machine learning tasks, RVFL can bridge the gap left by traditional neural networks.
Building on the previous discussion of MVL and the advantages of RVFL over conventional ANNs and ML models, we see a clear path forward in fusing these methods. The primary reason for this fusion is that the resulting model will be simple, scalable, efficient, and highly effective for DNA-binding protein prediction. Hence, we propose a novel multiview random vector functional link network (MvRVFL). MvRVFL adopts the framework of artificial neural networks combined with MVL, mirroring the structure of RVFL. To ensure the proposed model remains simple and effective, we incorporate two views at a time in the MvRVFL model. This approach strikes a balance between complexity and performance, leveraging complementary information from two views to enhance effectiveness while maintaining simplicity.
The following are the key features of the proposed MvRVFL model:
@@
    Integrating features from diverse perspectives enhances the proposed MvRVFL model’s capability to capture intricate patterns and relationships within the data, consequently enhancing the overall generalization performance of the model.
@@
    In the primary objective function of the proposed MvRVFL model, we encompass a coupling term, aiming to minimize a composite of errors originating from all views. By amalgamating the advantages of both early and late fusion, the model can assimilate information from all views during the training phase, yet still permits some flexibility to model the views distinctly.
@@
    The integration of multiple views in the proposed MvRVFL model helps to mitigate the impact of missing or noisy data in any single view (a common problem in biomedical datasets (Emmanuel et al., 2021)), thereby enhancing the model’s robustness and reliability.
@@
  • We propose a novel multiview random vector functional link (MvRVFL) network. The proposed MvRVFL leverages multiple distinct feature sets using the MVL and the simplicity cum effectiveness of RVFL to predict DBP.
  • We provide rigorous mathematical frameworks for MvRVFL, leveraging the RVFL topology. We utilized the Rademacher complexity theory to examine the consistency error bound and the generalization error bound of the proposed MvRVFL.
@@
  • Following (Guan et al., 2022), in this work, we partition each protein sequence into five equal-length segments, creating a total of 14 continuous or discontinuous regions. Each region is characterized using 63-dimensional features (7 + 21 + 35), leading to a comprehensive 882-dimensional (63 x 14) feature representation of the entire protein sequence. We extract five types of features from the sequences: NMBAC, MCD, PSSM-AB, PSSM-DWT, and PsePSSM. Experiments showcased that our method surpasses traditional models in effectively identifying DNA-binding proteins.
@@
  • Moreover, to test the generalization performance of the proposed model, we further tested it over the UCI, KEEL, AwA, and Corel5k datasets from various domains. The empirical results illustrate that the proposed MvRVFL model exhibits superior performance compared to numerous baseline models.
@@
The remaining structure of the paper is organized as follows. We discuss the literature on the baseline model in Section 2. Section 3, provides an overview of related work. Section 4 covers the method for extracting features from DNA-binding protein sequences. We present the detailed mathematical formulation of the proposed MvRVFL model in Section 5. Experimental results and analyses of proposed and existing models are discussed in Section 6. Finally, Section 7 presents the conclusions and potential future research directions.
2 Background
First, we explore the importance and function of DNA-binding proteins, which play a crucial role in various biological processes by interacting with DNA sequences. Next, we delve into the fundamentals of artificial neural networks (ANNs), a type of machine-learning model inspired by the human brain, and examine their architecture, learning mechanisms, and applications. Finally, we cover the concept of multi-view learning, a technique that integrates multiple sets of features or perspectives to improve the performance of the predictive model.
2.1 DNA binding protein
A DBP is a protein that can physically interact with DNA through its internal binding domain. As one of the most common intracellular proteins, DBPs play a crucial role in influencing genome function by participating in transcription, DNA replication, and DNA repair (Zimmermann et al., 2020). The active role of DBPs in cellular processes underscores their importance. As new proteins continue to be discovered in the post-genomic era, the identification of DBPs across various sequences has gained considerable attention. How can DBPs be identified among the multitude of newly discovered proteins? Experimental detection methods, which are resource-intensive, are not as cost-effective as computational approaches. As a result, recent years have seen the development of numerous computational predictive models for identifying DBPs. (Lin et al., 2011) introduced iDNA-Prot, which utilizes the pseudo amino acid composition (PseACC) (Fu et al., 2019) for feature extraction and applies a random forest (RF) classifier. Subsequently, Liu et al. developed three successive predictors: iDNA-Prot|dis (Liu et al., 2014), PseDNA-Pro (Liu et al., 2015c), and iDNAPro-PseAAC (Liu et al., 2015b). These predictors combined features from various extraction algorithms and used the integrated features as input for SVM to make predictions. Similarly, StackPDB predicts DNA-binding proteins through a three-step process: feature extraction, feature selection, and model construction are the key stages. StackPDB extracts features from protein sequences based on amino acid composition and evolutionary information. Evolutionary information is captured using the position-specific scoring matrix (PSSM), which is generated by the PSI-BLAST (Altschul et al., 1997b) program. In the StackPDB method, PsePSSM, PSSM-TPC, EDT, and RPT are employed for extracting features from PSSM. They subsequently used extreme gradient boosting combined with recursive feature elimination to select the most effective features. Finally, the selected optimal feature subset is input into a stacked ensemble classifier comprising XGBoost, SVM, and LightGBM. Previous studies (Ding et al., 2017; 2019) demonstrate that protein sequences can be characterized through different representations, including amino acid composition and PSSM. As fusion methods can combine information from multiple representations to improve model performance, various fusion techniques are employed in the identification of DBPs.
Multiple kernel learning (MKL) is a popular early fusion technique that focuses on learning the optimal weights for kernels. The optimal kernel is created by linearly combining multiple base kernels according to their respective weights. CKA-MKL (Qian et al., 2021) seeks to maximize the cosine similarity between the optimal kernel and the ideal kernel. Additionally, CKA-MKL includes a Laplacian term associated with the weights in the objective function to avoid extreme scenarios. However, CKA-MKL primarily emphasizes global kernel alignment and does not account for the differences between local samples. Therefore, HKAM-MKL (Zhao et al., 2022a) strives to maximize both local and global kernel alignment scores. Both CKA-MKL and HKAM-MKL utilize SVM as their classifier. In contrast, HSIC-MKL (Qian et al., 2022) aims to maximize the independence between trained samples and labels within the Reproducing Kernel Hilbert Space (RKHS). The optimal kernel is then used as input for a hypergraph-based Laplacian SVM, an extension of the standard SVM. In contrast, CKA-MKL focuses solely on global alignment. Moreover, HKAM-MKL takes into account both global and local aspects. As a result, HKAM-MKL outperforms CKA-MKL in predicting DNA-binding proteins. Unlike the previously mentioned MKL methods, MLapSVM-LBS (Sun et al., 2022) integrates multiple types of information throughout the training process. It uses the multiple local behavior similarity graph as a regularization term. Given that the objective function of MLapSVM-LBS is non-convex, an alternating algorithm is applied. The key advantage of MLapSVM-LBS is its ability to fuse various sources of information during training while providing flexibility to model different views distinctly.
2.2 Artificial neural networks
Artificial neural networks (ANNs) are a category of machine learning models inspired by the neural structure of the human brain. ANNs comprise interconnected nodes, or neurons, which employ mathematical operations to process and transmit information. ANNs are engineered to identify patterns and correlations within data, leveraging this acquired knowledge to make predictions. ANNs have showcased efficacy across diverse domains, including brain age prediction (Tanveer et al., 2023), fault diagnosis of drilling pumps (Guo et al., 2024), detection of sickle cell disease (Goswami et al., 2024), rainfall forecasting (Luk et al., 2001), diagnosis of Alzheimer’s disease (Tanveer et al., 2024; Ganaie et al., 2024), and so on. Conventional ANNs rely on gradient descent (GD) based iterative methods, which present several inherent challenges in parameter calculation. These include a tendency to converge to local rather than global optima, heightened sensitivity to the choice of learning rate and initial parameters, and a sluggish convergence rate.
To circumvent the limitations of GD-based neural networks, randomized neural network (RNN) (Schmidt et al., 1992) is proposed. In RNNs, certain network parameters remain fixed, with only the parameters of the output layer being computed using a closed-form solution throughout the training phase (Suganthan & Katuwal, 2021). The random vector functional link (RVFL) neural network (Pao et al., 1994; Malik et al., 2023) is a shallow feed-forward RNN distinguished by randomly initialized hidden layer parameters, which remain constant during the training process. RVFL stands out among other RNNs because of its direct connections between output and input layers. These direct links serve as a type of implicit regularization (Zhang & Suganthan, 2016; Sajid et al., 2024) within RVFL, contributing to enhanced learning capabilities. Through methods such as the least-squares technique or Pseudo-inverse, RVFL provides a closed-form solution for optimizing output parameters. This characteristic leads to efficient learning with fewer adjustable parameters. For more details, interested readers can refer to the comprehensive review paper of RVFL (Malik et al., 2023).
2.3 Multiview learning
Multiview learning (MVL), as a prominent research field, showcases the ability to substantially improve generalization performance across diverse learning tasks by integrating multiple feature sets that encompass complementary information (Zhao et al., 2017b; Xu et al., 2017; Tang et al., 2020). MVL emerges in response to the common occurrence of diverse types of data in practical scenarios. Consider an image, which can be characterized by its color or texture features, and a person, who can be identified through facial characteristics or fingerprints. In real-world situations, samples from different perspectives might exist in separate spaces or display notably different distributions because of the significant variation between views. However, the naive methods address this data type by employing a cascade strategy (Xu et al., 2017). This entails transforming the multiview data into new single-view data by consolidating the heterogeneous feature space into a homogeneous feature space. However, the cascading strategy overlooks the unique statistical properties of each view and is plagued by the curse of dimensionality. MVL techniques (Tang et al., 2019; Huang et al., 2020) are applied across various tasks, including transfer learning (Zhao et al., 2017a), clustering (Wen et al., 2018; 2020), dimensionality reduction (Wang & Zhang, 2013; Hu et al., 2019), and classification (Sun et al., 2018). SVM-2K, a two-view SVM learning model that combines SVMs with the kernel canonical correlation analysis (KCCA) distance minimization version, was first introduced by (Farquhar et al., 2005). Using the consensus principle, this method makes use of two points of view. Multiview twin SVM (MvTSVM) (Xie & Sun, 2015) represents the initial endeavor to integrate a best-fitting hyperplane classifier with MVL. In recent years, various variants of MvTSVM have been introduced such as multiview one-class SVM method with LUPI (MOCPIL) (Xiao et al., 2024), multiview large margin distribution machine (MVLDM) (Hu et al., 2024), multiview restricted kernel machine (MVRKM) (Houthuys & Suykens, 2021), and multiview robust double-sided TSVM (MvRDTSVM) (Ye et al., 2021).
3 Related Work
This section begins with establishing notations and then reviews the mathematical formulation along with the solution of the RVFL network.
3.1 Notations

### v2

@@ Abstract
The identification of DNA-binding proteins (DBPs) is essential due to their significant impact on various biological activities. Understanding the mechanisms underlying protein-DNA interactions is essential for elucidating various life activities. In recent years, machine learning-based models have been prominently utilized for DBP prediction. In this paper, to predict DBPs, we propose a novel framework termed a multiview random vector functional link (MvRVFL) network, which fuses neural network architecture with multiview learning. The MvRVFL model integrates both late and early fusion advantages, enabling separate regularization parameters for each view, while utilizing a closed-form solution for efficiently determining unknown parameters. The primal objective function incorporates a coupling term aimed at minimizing a composite of errors stemming from all views. From each of the three protein views of the DBP datasets, we extract five features. These features are then fused together by incorporating a hidden feature during the model training process. The performance of the proposed MvRVFL model on the DBP dataset surpasses that of baseline models, demonstrating its superior effectiveness. We further validate the practicality of the proposed model across diverse benchmark datasets, and both theoretical analysis and empirical results consistently demonstrate its superior generalization performance over baseline models. The source code of the proposed MvRVFL model is available at https://github.com/mtanveer1/MvRVFL.
keywords:
@@
Multiview learning, Random vector functional link network, Extreme learning machine, Support vector machine, DNA-binding protein.
journal: Elsevier
[…]
[inst1]organization=Department of Mathematics, Indian Institute of Technology Indore,addressline=Simrol, city=Indore, postcode=453552, state=Madhya Pradesh, country=India
1 Introduction and Motivation
@@
A DNA-binding protein (DBP) is a protein that has the ability to interact with DNA [33]. The DBP plays a crucial role in numerous vital biological processes, including transcriptional regulation, DNA replication and repair, cellular development, and chromatin organization [45]. Therefore, precise prediction of DBPs is of considerable significance for proteome annotation as well as endeavors in synthetic biology. To identify DBPs, extensive work has been carried out in wet lab settings, including techniques such as X-Ray crystallography [8], genetic analysis [19], and chromatin immunoprecipitation on microarrays [6]. While results obtained from wet-lab methods are likely the most reliable, it’s important to note that this approach is also associated with significantly high time and labor costs. In contrast to wet-lab methods, computational approaches can substantially decrease resource requirements and expedite the identification of DBPs. Furthermore, in the postgenomic era, there’s an escalating demand for the development of efficient and swift computational methods for accurately identifying DBPs, thereby underscoring their increasing significance within the bioinformatics domain.
@@
Computational techniques are widely used for identifying DBPs [50]. These techniques are generally divided into two primary types: structure-based methods and sequence-based methods. Structure-based methods utilize information derived from the protein’s three-dimensional structure. Initially, structure-based methods dominated the field of DBP prediction, with models such as LBi-DBP [74], StackDPP [1], DBPboost [58], and ULDNA [83]. Effectively encoding a protein sequence into a vector format is considered both a fundamental and difficult aspect of sequence-based approaches [75]. DNABinder [52] employs evolutionary information encoded in a position-specific scoring matrix (PSSM), generated through the PSI-BLAST multiple sequence alignment tool [2]. A support vector machine (SVM) is utilized for classification, taking these features as input, representing the initial application of this method in identifying DNA-binding proteins (DBPs). In iDNA-Prot [9], pseudo amino acid composition (PseAAC) features [11], derived from protein sequences through the grey model, are combined with random forest (RF) for identifying DNA-binding proteins (DBPs). iDNAProt-ES [11] utilizes evolutionary data like PSSM alongside structural information predicted by SPIDER2 [72] to characterize the features of a given protein. These features are trained to identify DBPs using SVM with a linear kernel. DBP datasets, derived from various sources or views, offer complementary information essential for improving task performance. Each view contributes unique perspectives and features that, when combined, enhance the overall understanding and efficiency of the model. Nevertheless, existing multiview learning approaches often necessitate complex architectures or extensive preprocessing to manage these diverse views effectively [28, 51]. The growing complexity of data from different sources poses significant challenges in extracting meaningful insights and achieving high performance. In recent years, the integration of multiview information has become increasingly popular in the domain of bioinformatics [85, 82]. The FKRR-MVSF model [85], which utilizes multi-view sequence features and fuzzy kernel ridge regression, is introduced for the identification of DNA-binding proteins. In FKRR-MVSF, the initial step entails deriving multi-view sequence features from protein sequences, followed by the application of a multiple kernel learning (MKL) approach to combine these varied features. In MSFBinder [40], a stacking framework is introduced to predict DNA-binding proteins (DBPs) by merging features from various perspectives. The multi-view hypergraph restricted kernel machine (MV-H-RKM) [25] model is introduced to derive various features from protein sequences. The multi-view features are linked through a shared latent representation, and multi-hypergraph regularization is employed to integrate them while preserving the structural coherence between the original features and the latent space.
Despite the extensive use of hyperplane-based classifiers like SVMs and their variants in predicting DNA-binding proteins, however, DBP prediction remains a challenging task due to the complex, heterogeneous, and nonlinear nature of protein sequences, which encapsulate diverse evolutionary, structural, and physicochemical properties [23]. SVM-based models and manual identification are both time-consuming and non-scalable, highlighting the growing need for efficient machine learning–based predictive frameworks [59].
Among efficient learning paradigms, the Random Vector Functional Link (RVFL) [46, 43] network has been widely recognized for its fast training speed and closed-form learning capability [54, 55]. Despite these advantages, conventional RVFL architectures fail to fully leverage complementary information from multiple heterogeneous feature representations, as evidenced by our experimental observations [53]. This shortcoming arises because standard RVFLs perform shallow transformations of input data, limiting their capacity to capture intricate cross-feature dependencies, particularly when integrating PSSM-based, evolutionary, and physicochemical descriptors.
In the context of DBP prediction, there exists a pressing need for an RVFL-based model that can effectively exploit the complementary nature of multiple biological feature views while preserving computational efficiency. To address this gap, we introduce the Multiview RVFL (MvRVFL) network, an advanced RVFL formulation designed to integrate diverse biological features through a unified multiview learning framework. Unlike deep learning models that require extensive parameter tuning and computational resources, MvRVFL preserves the analytical training property of classical RVFL, yielding a closed-form solution with low memory and runtime overhead.
The main characteristics of the proposed MvRVFL model are as follows:
@@
    To the best of our knowledge, this is the first time to integrate multiview learning principles into the RVFL network. While conventional RVFL models are known for their simplicity, scalability, and efficiency compared to traditional ANNs and ML methods, they have not been extended to a principled multiview optimization setting. The proposed MvRVFL formulation enables simultaneous exploitation of complementary information through both early and late fusion across distinct biological feature views.
@@
    We introduce a novel optimization framework for MvRVFL that incorporates a coupling term into the objective function. This design allows the model to minimize a composite of errors originating from all views while maintaining a balance between shared and view-specific learning. The formulation thus provides a mathematically grounded mechanism for effective multiview unification.
@@
    The proposed model unifies early fusion (which captures inter-view correlations at the input level) and late fusion (which aggregates decision-level outputs from each view). This hybrid fusion design enables MvRVFL to capture shared structures while preserving individual view information. The ablation studies empirically confirm that this combined fusion approach consistently outperforms models employing only early or late fusion, thereby validating its effectiveness (See section 6.4).
  4. 4.
    The proposed MvRVFL framework is specifically designed with the goal of enhancing DBP prediction. By leveraging heterogeneous protein descriptors, such as PSSM-based, evolutionary, and physicochemical features, as distinct views, MvRVFL effectively captures complex discriminative patterns.
@@
    We propose a novel MvRVFL network that leverages multiple distinct feature sets using the MVL and the simplicity cum effectiveness of RVFL to predict DBP.
  • 2.
    The mathematical foundations of the MvRVFL model are established by deriving its consistency and generalization error bounds using Rademacher complexity analysis.
  • 3.
@@
    Following the method outlined in [25], we partition each protein sequence into five segments of equal size, resulting in 14 distinct continuous and discontinuous regions. Each region is characterized using 63-dimensional features (7 + 21 + 35), leading to a comprehensive 882-dimensional (63 x 14) feature representation of the entire protein sequence. We extract five types of features from the sequences: NMBAC, PSSM-DWT, MCD, PSSM-AB, and PsePSSM. Experiments showcased that our method outperforms traditional models in effectively identifying DNA-binding proteins.
  • 4.
@@
    Moreover, to assess the generalization capability of the proposed model, we perform evaluations using the UCI, KEEL, Corel5k, and AwA datasets, which span various domains. The results from the experiments show that the MvRVFL model surpasses various baseline models in performance.
@@
The remainder of the paper is organized as follows: A offers a literature review on the baseline model, and B provides an overview of related work. Section 2 introduces the notations and mathematical symbols that are consistently used throughout the paper. Section 3 covers the method for extracting features from DNA-binding protein sequences. We present the detailed formulation of the proposed MvRVFL model in Section 4. Section 5 explores the ability of the proposed MvRVFL model to generalize effectively. In Section 6, we detail the experimental findings and offer a comparative assessment between the proposed approach and established baseline models. Finally, Section 7 presents the conclusions and potential future research directions.
2 Notations